OpenStudy (anonymous):

Find x - 3y if x = 3 and y = -2

OpenStudy (anonymous):

Just plug in x and y into x and y in the equation o_O

OpenStudy (anonymous):

x-3y =(3)-3(-2) =3+6 =9

OpenStudy (anonymous):

-.-